OpenStudy (anonymous):

a^2+8a+15/a^2+7a+12 * 3a+12/a^2-25 someone please help

mathslover (mathslover):

is it like this : \[\huge{\frac{a^2+8a+15}{a^2+7a+12}*\frac{3a+12}{a^2-25}}\]

OpenStudy (anonymous):

yes

mathslover (mathslover):

ok then : \[\huge{\frac{a^2+8a+15}{a^2+7a+12}}\] i am simplifying this term first : \[\huge{\frac{a^2+5a+3a+15}{a^2+4a+3a+12}}\] \[\huge{\frac{(a+5)(a+3)}{(a+3)(a+4)}}\]

mathslover (mathslover):

similarly simplify second term . what do you get?

OpenStudy (anonymous):

8a^2/7a^2 ?

mathslover (mathslover):

wait let me check

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

I actually have three choices. they are -3/a+4 3(a+5)/a^2-25 3/a-5 @mathslover

mathslover (mathslover):

3/a-5 is correct

OpenStudy (anonymous):

alright thank you so much